OpenStudy (anonymous):

Simplify 2x - 8x - 6y + 9 -14 + x + 11y. 5x - 5y - 5 -5x + 5y - 5 5x + 5y + 5 -5x - 5y + 5

OpenStudy (unknownrandom):

Combine like terms it will be the second one. -5x+5y-5

OpenStudy (anonymous):

can u explain how u got that

OpenStudy (unknownrandom):

You have to treat each letter as it's own problem.

OpenStudy (the_fizicx99):

2x - 8x - 6y + 9 -14 + x + 11y -6x -6y + 9 -14 + x + 11y I use parenthesis to separate them, (-6x) + (x) = -5x -6y + 11y = 5y 9 - 14 = -5 -5x + 5y - 5

OpenStudy (unknownrandom):

So first x -8x+(2x+x) -8x+3x -5x

OpenStudy (the_fizicx99):

You gave him the medal because he said its A, B, C or D?!?! Done~~

OpenStudy (usukidoll):

huh?

OpenStudy (unknownrandom):

You do that for each letter. Do you understand how we got that @parisgirl1515?

OpenStudy (the_fizicx99):

Who did you huh at usuki :#

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

u @tHe_FiZiCx99 about medals

OpenStudy (the_fizicx99):

I don't get it ._. did she give him the medal because he said it was the second? Like shows she only wants answers in my opinion.. @UsukiDoll

OpenStudy (anonymous):

no I check my answers

OpenStudy (the_fizicx99):

Idc about medals, its virtual so no benefit to me.. I joined to help lol, its worthless to meh :3

OpenStudy (usukidoll):

I gave the medal to unknown

OpenStudy (anonymous):

okay I will give u a stupid medal.

OpenStudy (usukidoll):

-.- excuse me...wanna trade problems with me?

OpenStudy (usukidoll):

my material is much harder than yours ;)

OpenStudy (anonymous):

okay bye guys

OpenStudy (the_fizicx99):

I don't want your medal =.= * walks along usuki*

OpenStudy (usukidoll):

rofl the

OpenStudy (anonymous):

Im just checking my hw thats it

OpenStudy (usukidoll):

ok nothing wrong with that ^^

OpenStudy (anonymous):

okay

OpenStudy (the_fizicx99):

Usuki :*

OpenStudy (usukidoll):

yes?

OpenStudy (the_fizicx99):

Bestie